A Bibliography of Publications in Biometrics: 1950–1959

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

25 June 2020 Version 1.00

Title word cross-reference

2 × 2 [47, 46]. 3 × 6 [61]. 4^{n} [62]. χ^{2} [44, 12]. F [48]. k_{4} [71]. r × c [12]. X^{2} [23, 45].

113 [50].

8 [100].

abstract [99]. adjusted [75]. age [57]. age-specific [57]. among [6]. analyses [82, 71]. Analysis [80, 35, 18, 41, 70, 53, 13, 5, 65, 68, 38, 56, 61, 79, 28, 69, 85, 90, 3, 81]. Answer [50]. application [76, 9]. Applications [71, 65]. applied [84]. arranged [61]. Asymptotic [10, 67, 89]. Automatic [23].

balance [74]. balanced [79, 8, 18, 16, 41]. between [15, 64]. beyond [65]. binomial [29, 30, 38, 84]. Bioassay [78]. bioassays [73]. biological [29]. biologie [9]. biology [33]. Biometrics [100]. block [15, 13, 65, 79, 40, 8, 18, 91, 16, 74, 52, 27, 41, 81]. blocks [28, 69]. both [72].

calculating [23]. calculation [12]. calculator [19, 39]. cas [11]. case [11]. causal [90]. censuses [21, 100]. certain [11]. certains [11]. Chain [40, 27]. Chapman [100]. class [91]. classes [59]. classifications [82]. coefficient [58, 94]. Combining [96]. comments [93]. common [44]. comparison [97]. comparisons [47, 70, 49]. complex [68]. component [5, 71]. Components [7, 32]. composite [98]. Compound [3]. computation [19]. computers [56]. computing [39]. concept [58]. concepts [14]. Confidence [73, 55, 54, 89]. connected [83]. constraints [68]. constructing [88]. contagious [34, 51]. contingency [45, 12, 23]. Contribution [93]. Contributions [54, 49]. control [97]. correlated [75]. correlation [24, 94]. covariance [80, 82, 32, 95, 85, 81]. curve [67, 57]. cut [45].

D [100]. dans [11]. data [29]. Design [28, 35, 52, 27]. designs [15, 65, 2, 98, 79, 40, 8, 18, 91, 16, 74, 41, 81]. desk [19, 39]. d’estimation [11]. difference [64]. differences [15]. different [59]. distribution [29, 51, 36, 63, 89, 84]. distribution-free [89]. distributions [34, 37, 83, 43, 14, 22]. divisible [79]. double [11]. dual [16].

échantillonnage [11]. effect [70]. effects [92]. efficient [30]. elimination [40]. Erratum [100]. error [1, 72, 52]. estimate [77]. estimates [93]. estimating [89]. Estimation [32, 59, 1, 11, 54, 14, 22, 57]. estimators [96]. exact [45]. Examples [79]. existing [65]. experiment [68, 61]. experimental [83]. experiments [62, 3]. Extension [66, 75]. Extensions [60].

factorial [47, 62]. factorials [79, 99]. Fisher [37]. Fitting [29, 51, 86, 30, 72, 67]. flavor [70]. formulas [45]. free [89]. French [11]. function [77]. functions [37]. Fundamental [14]. Further [49].

G [100]. general [2, 56, 14]. generalization [34]. genetic [55, 95, 94, 71]. genetics [58]. group [66, 79, 75]. grouped [93]. growth [31, 42].

having [41]. heritability [55, 89]. heterogeneity [40].

II [22]. illustrated [18]. impact [58]. Incomplete [74, 65, 79, 8, 18, 91, 16, 52, 41, 81]. independent [88]. indices [87]. infection [57]. information [13, 74]. inter [13, 74, 52]. inter- [52]. inter-block [13, 74]. interval [54]. intervals [55]. intra [79, 52]. intra-block [79, 52]. Inverse [21, 100]. involving [68]. IV [93].

Kronecker [76].

Latin [61]. Latinized [17]. lattice [2, 35]. lattices [13, 1, 17, 8, 18, 26]. learning [92]. Lee [37]. likelihood [77, 86, 84]. limits [73, 89]. line [72, 24]. linear [6]. linked [28, 69]. Little [99]. logistic [77, 86].

machine [23]. macronuclear [42]. Matched [64]. mathematical [4]. mathématique [9]. Matrices [53, 76]. maximum [77, 86, 84]. mean [25]. means [15, 66, 75]. measuring [73, 70]. medical [3]. method [92, 72, 67, 88, 89, 23, 84]. methods [44, 14]. Misclassification [46]. missing [15, 60, 41]. mixed [99]. model [42]. moments [39]. Multi [22]. Multi-stimulus [22]. multinomial [83]. Multiple [48, 21, 76, 66, 75, 100, 97]. multivariate [85, 3].

nature [80]. negative [29, 30]. Neyman [34, 51]. normal [37]. Note [30, 72, 62, 65, 43, 26]. number [59]. numbers [66].

observations [93]. operations [39]. organic [24]. orthogonal [88]. Outline [4].

paired [47, 70, 49]. pairs [64]. parabola [78]. paramecium [42]. parameter [51]. partially [79, 8, 18, 41]. Partition [83, 45]. path [58]. paths [90]. Pearson [37]. peck [4]. pieces [99]. plan [56]. Pleiotropism [95]. plot [60]. Poisson [20, 43, 36, 63, 84]. polynomials [88]. Population [31, 59, 58]. power [2]. precision [73]. presence [92]. present [5]. prime [2]. prime-power [2]. problèmes [11]. problems [11]. processes [33]. product [76]. products [19]. programming [56]. proportions [64]. purpose [56].

quadratic [98]. quantal [53]. quasi [61]. quasi-Latin [61]. Query [50].

randomized [15]. range [48, 66, 75]. rank [65, 97]. rankit [70]. rates [57]. ratios [55]. recovery [74]. Rectangular [8, 13, 1, 17, 35, 18, 26]. regression [76, 67, 10, 7, 90]. relation [6]. relative [52, 57]. replicate [91]. replications [66]. response [98, 14, 22]. response-time [14, 22]. Restricted [87]. right [4]. row [39, 74].

sample [21, 63, 100]. sampling [11, 94]. scope [65]. selection [87, 42]. sequential [64, 21, 100]. series [62]. setting [89]. sexes [31]. Short [45]. Short-cut [45]. significance [64, 63]. simple [93, 18, 67, 88]. Simplified [69]. simultaneous [54]. singly [37, 28, 69]. size [52]. sizes [13]. small [63]. Smallest [98]. Some [44, 63, 16, 41, 93]. spaced [88]. specific [57]. specifying [55]. sprays [70]. square [61, 18]. squares [19]. statistic [71]. statistical [68]. statistique [9]. status [5]. stimulus [22]. stochastic [42, 33]. straight [72]. strengthening [44]. study [93]. subject [72]. Suggested [39]. sum [97]. sums [19]. surfaces [98]. symmetry [3].

table [12]. Tables [77, 37, 46, 65, 45]. techniques [60]. test [97]. Testing [6, 92, 89]. tests [44, 48, 66, 75, 63, 3]. theory [2, 49, 4]. there [15]. time [14, 22]. transformations [38]. treatment [92]. treatments [97]. trials [64]. triangular [28]. triple [35]. truncated [37, 20, 43, 36]. two [15, 51, 40, 91]. two-way [40]. Type [51].

unbalanced [82]. unbiased [96]. unequal [13, 66]. unequally [88]. unintentional [68]. uniqueness [24]. units [42]. uses [80]. using [13].

values [15]. variable [88]. variables [72]. Variance [82, 25, 5, 38, 55, 56, 32, 95, 94, 71]. Variances [15, 6]. various [38]. vectors [83]. versus [97].

Wald [72]. way [40]. weighted [25].

References

Grundy:1950:EER

[1] P. M. Grundy. The estimation of error in rectangular lattices. Biometrics, 6(??):25–33, 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Federer:1950:GTP

[2] W. T. Federer. The general theory of prime-power lattice designs. Biometrics, 6(??):34–58, 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Votaw:1950:CST

[3] D. F. Votaw, Jr., A. W. Kimball, and J. A. Rafferty. Compound symmetry tests in the multivariate analysis of medical experiments. Biometrics, 6 (??):259–281, 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rapoport:1950:OMT

[4] Anatol Rapoport. Outline of a mathematical theory of peck right. Biometrics, 6(??):330–341, 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Crump:1951:PSV

[5] S. Lee Crump. The present status of variance component analysis. Biometrics, 7(??):1–16, 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cochran:1951:TLR

[6] W. G. Cochran. Testing a linear relation among variances. Biometrics, 7 (??):17–32, 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tukey:1951:CR

[7] John W. Tukey. Components in regression. Biometrics, 7(??):33–69, ???? 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nair:1951:RLP

[8] K. R. Nair. Rectangular lattices and partially balanced incomplete block designs. Biometrics, 7(??):145–154, 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Frechet:1951:ASM

[9] Maurice Fréchet. Sur une application de la statistique mathématique à la biologie. Biometrics, 7(??):180–184, 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stevens:1951:AR

[10] W. L. Stevens. Asymptotic regression. Biometrics, 7(??):247–267, 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lamotte:1951:CPE

[11] M. Lamotte and M. Schutzenberger. Sur certains problèmes d’estimation dans les cas de double échantillonnage. (French) [On certain estimation problems in the case of double sampling]. Biometrics, 7(??):275–282, 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Leslie:1951:CCT

[12] P. H. Leslie. The calculation of χ^{2} for an r × c contingency table. Biometrics, 7(??):283–286, 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Clarke:1951:ARL

[13] Pamela M. Clarke. An analysis of rectangular lattices with unequal block sizes, using inter-block information. Biometrics, 7(??):287–294, 1951. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sampford:1952:ERTa

[14] M. R. Sampford. The estimation of response-time distributions. I. Fundamental concepts and general methods. Biometrics, 8(??):13–32, 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baten:1952:VDB

[15] William Dowell Baten. Variances of differences between means when there are two missing values in randomized block designs. Biometrics, 8(??):42– 50, 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shrikhande:1952:DSB

[16] S. S. Shrikhande. On the dual of some balanced incomplete block designs. Biometrics, 8(??):66–72, 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Harshbarger:1952:LRL

[17] Boyd Harshbarger and Lyle L. Davis. Latinized rectangular lattices. Biometrics, 8(??):73–84, 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nair:1952:APB

[18] K. R. Nair. Analysis of partially balanced incomplete block designs illustrated on the simple square and rectangular lattices. Biometrics, 8(??): 122–155, 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hammersley:1952:CSS

[19] J. M. Hammersley. The computation of sums of squares and products on a desk calculator. Biometrics, 8(??):156–168, 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

David:1952:TP

[20] F. N. David and N. I. Johnson. The truncated Poisson. Biometrics, 8(??): 275–285, 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chapman:1952:IMS

[21] Douglas G. Chapman. Inverse, multiple and sequential sample censuses. Biometrics, 8(??):286–306, 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See erratum [100].

Sampford:1952:ERTb

[22] M. R. Sampford. The estimation of response-time distributions. II. Multi-stimulus distributions. Biometrics, 8(??):307–369, 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Skory:1952:AMM

[23] John Skory. Automatic machine method of calculating contingency X^{2}. Biometrics, 8(??):380–382, 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kruskal:1953:ULO

[24] William H. Kruskal. On the uniqueness of the line of organic correlation. Biometrics, 9(??):47–58, 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Meier:1953:VWM

[25] Paul Meier. Variance of a weighted mean. Biometrics, 9(??):59–73, 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nair:1953:NRL

[26] K. R. Nair. A note on rectangular lattices. Biometrics, 9(??):101–106, 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Youden:1953:CBD

[27] W. J. Youden and W. S. Connor. The chain block design. Biometrics, 9(??):127–140, 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nair:1953:DAT

[28] K. R. Nair. Design and analysis of triangular singly linked blocks. Biometrics, 9(??):141–156, 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bliss:1953:FNB

[29] C. I. Bliss. Fitting the negative binomial distribution to biological data. Biometrics, 9(??):176–196, 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fisher:1953:NEF

[30] R. A. Fisher. Note on the efficient fitting of the negative binomial. Biometrics, 9(??):197–200, 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goodman:1953:PGS

[31] Leo A. Goodman. Population growth of the sexes. Biometrics, 9(??): 212–225, 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Henderson:1953:EVC

[32] C. R. Henderson. Estimation of variance and covariance components. Biometrics, 9(??):226–252, 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reid:1953:SPB

[33] A. T. Reid. On stochastic processes in biology. Biometrics, 9(??):275–289, 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Beall:1953:GNC

[34] Geoffrey Beall and Richard R. Rescia. A generalization of Neyman’s contagious distributions. Biometrics, 9(??):354–386, 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Murty:1953:ATR

[35] V. N. Murty. Analysis of a triple rectangular lattice design. Biometrics, 9(??):422–424, 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Plackett:1953:TPD

[36] R. L. Plackett. The truncated Poisson distribution. Biometrics, 9(??): 485–488, 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cohen:1953:TPL

[37] A. C. Cohen, Jr. and John Woodward. Tables of Pearson–Lee–Fisher functions of singly truncated normal distributions. Biometrics, 9(??):489–497, 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fisher:1954:AVV

[38] Ronald Fisher. The analysis of variance with various binomial transformations. Biometrics, 10(??):130–139; discussion, 140–151, 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hemphill:1954:SDC

[39] F. M. Hemphill. Suggested desk calculator operations for computing moments by the row. Biometrics, 10(??):152–154, 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mandel:1954:CBD

[40] John Mandel. Chain block designs with two-way elimination of heterogeneity. Biometrics, 10(??):251–272, 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zelen:1954:ASP

[41] Marvin Zelen. Analysis for some partially balanced incomplete block designs having a missing block. Biometrics, 10(??):273–281, 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kimball:1954:SMS

[42] A. W. Kimball and A. S. Householder. A stochastic model for the selection of macronuclear units in paramecium growth. Biometrics, 10(3):361–374, September 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001591.

Moore:1954:NTP

[43] P. G. Moore. A note on truncated Poisson distributions. Biometrics, 10 (??):402–406, 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cochran:1954:SMS

[44] William G. Cochran. Some methods for strengthening the common χ^{2} tests. Biometrics, 10(??):417–451, 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kimball:1954:SCF

[45] A. W. Kimball. Short-cut formulas for the exact partition of x^{2} in contingency tables. Biometrics, 10(??):452–458, 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bross:1954:MT

[46] Irwin Bross. Misclassification in 2 × 2 tables. Biometrics, 10(??):478–486, 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Abelson:1954:FPC

[47] Robert M. Abelson and Ralph Allan Bradley. A 2×2 factorial with paired comparisons. Biometrics, 10(??):487–502, 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Duncan:1955:MRM

[48] David B. Duncan. Multiple range and multiple F tests. Biometrics, 11 (??):1–42, 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kendall:1955:FCT

[49] M. G. Kendall. Further contributions to the theory of paired comparisons. Biometrics, 11(??):43–62, 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tukey:1955:AQ

[50] John W. Tukey. Answer to Query 113. Biometrics, 11(??):111–113, ???? 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Douglas:1955:FNT

[51] J. B. Douglas. Fitting the Neyman Type A (two parameter) contagious distribution. Biometrics, 11(??):149–173, 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thompson:1955:RSI

[52] W. A. Thompson, Jr. The relative size of the inter- and intra-block error in an incomplete block design. Biometrics, 11(??):406–426, 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Claringbold:1955:MQA

[53] P. J. Claringbold. Matrices in quantal analysis. Biometrics, 11(??):481– 501, 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ramachandran:1956:CSC

[54] K. V. Ramachandran. Contributions to simultaneous confidence interval estimation. Biometrics, 12(??):51–56, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Graybill:1956:CIV

[55] Franklin A. Graybill, Frank Martin, and George Godfrey. Confidence intervals for variance ratios specifying genetic heritability. Biometrics, 12 (??):99–109, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hartley:1956:PPA

[56] H. O. Hartley. A plan for programming analysis of variance for general purpose computers. Biometrics, 12(??):110–122, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Whittle:1956:EAS

[57] P. Whittle. The estimation of age-specific infection rates from a curve of relative infection. Biometrics, 12(??):154–162, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:1956:CPC

[58] C. C. Li. The concept of path coefficient and its impact on population genetics. Biometrics, 12(??):190–210, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lewontin:1956:END

[59] R. C. Lewontin and Timothy Prout. Estimation of the number of different classes in a population. Biometrics, 12(??):211–223, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thompson:1956:EMP

[60] H. R. Thompson. Extensions to missing plot techniques. Biometrics, 12 (??):241–244, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hodnett:1956:AEA

[61] G. E. Hodnett. The analysis of a 3 × 6 experiment arranged in a quasi-Latin square. Biometrics, 12(??):245–258, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Claringbold:1956:NSF

[62] P. J. Claringbold. A note on the 4^{n} series of factorial experiments. Biometrics, 12(??):259–263, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rao:1956:SSS

[63] C. Radhakrishna Rao and I. M. Chakravarti. Some small sample tests of significance for a Poisson distribution. Biometrics, 12(??):264–282, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Billewicz:1956:MPS

[64] W. Z. Billewicz. Matched pairs in sequential trials for significance of a difference between proportions. Biometrics, 12(??):283–300, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dykstra:1956:NRA

[65] Otto Dykstra, Jr. A note on the rank analysis of incomplete block designs — applications beyond the scope of existing tables. Biometrics, 12(??): 301–306, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kramer:1956:EMR

[66] Clyde Young Kramer. Extension of multiple range tests to group means with unequal numbers of replications. Biometrics, 12(??):307–310, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Patterson:1956:SMF

[67] H. D. Patterson. A simple method for fitting an asymptotic regression curve. Biometrics, 12(??):323–329, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Finney:1956:SAC

[68] D. J. Finney and F. W. Cope. The statistical analysis of a complex experiment involving unintentional constraints. Biometrics, 12(??):345–368, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nair:1956:SAS

[69] K. R. Nair. Simplified analysis of singly linked blocks. Biometrics, 12(??): 369–380, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bliss:1956:RAP

[70] C. I. Bliss, Mary L. Greenwood, and Edna Sakamoto White. A rankit analysis of paired comparisons for measuring the effect of sprays on flavor. Biometrics, 12(??):381–403, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Robson:1956:ASG

[71] D. S. Robson. Applications of the k_{4} statistic to genetic variance component analyses. Biometrics, 12(??):433–444, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Keeping:1956:NWM

[72] E. S. Keeping. Note on Wald’s method of fitting a straight line when both variables are subject to error. Biometrics, 12(??):445–448, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bliss:1956:CLM

[73] C. I. Bliss. Confidence limits for measuring the precision of bioassays. Biometrics, 12(??):491–526, 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Taylor:1957:IBD

[74] W. B. Taylor. Incomplete block designs with row balance and recovery of inter-block information. Biometrics, 13(??):1–12, 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kramer:1957:EMR

[75] Clyde Young Kramer. Extension of multiple range tests to group correlated adjusted means. Biometrics, 13(??):13–18, 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cornish:1957:AKP

[76] E. A. Cornish. An application of the Kronecker product of matrices in multiple regression. Biometrics, 13(??):19–27, 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berkson:1957:TML

[77] Joseph Berkson. Tables for the maximum likelihood estimate of the logistic function. Biometrics, 13(??):28–34, 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bliss:1957:BP

[78] C. I. Bliss. Bioassay from a parabola. Biometrics, 13(??):35–50, 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kramer:1957:EIB

[79] Clyde Young Kramer and Ralph Allan Bradley. Examples of intra-block analysis for factorials in group divisible, partially balanced, incomplete block designs. Biometrics, 13(??):197–224, 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cochran:1957:ACN

[80] William G. Cochran. Analysis of covariance: Its nature and uses. Biometrics, 13(??):261–281, 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zelen:1957:ACI

[81] Marvin Zelen. The analysis of covariance for incomplete block designs. Biometrics, 13(??):309–332, 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Federer:1957:VCA

[82] Walter T. Federer. Variance and covariance analyses for unbalanced classifications. Biometrics, 13(??):333–362, 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Corsten:1957:PEV

[83] L. C. Corsten. Partition of experimental vectors connected with multinomial distributions. Biometrics, 13(??):451–484, 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sprott:1958:MML

[84] D. A. Sprott. The method of maximum likelihood applied to the Poisson binomial distribution. Biometrics, 14(??):97–106, 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Smith:1958:MAC

[85] H. Fairfield Smith. A multivariate analysis of covariance. Biometrics, 14 (??):107–127, 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hodges:1958:FLM

[86] J. L. Hodges, Jr. Fitting the logistic by maximum likelihood. Biometrics, 14(??):453–461, 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kempthorne:1959:RSI

[87] Oscar Kempthorne and Arne W. Nordskog. Restricted selection indices. Biometrics, 15(??):10–19, 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Robson:1959:SMC

[88] D. S. Robson. A simple method for constructing orthogonal polynomials when the independent variable is unequally spaced. Biometrics, 15(??): 187–191, 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schwartz:1959:DFA

[89] Lorraine Schwartz and Stanley Wearden. A distribution-free asymptotic method of estimating, testing, and setting confidence limits for heritability. Biometrics, 15(??):227–235, 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Turner:1959:RAC

[90] Malcolm E. Turner and Charles D. Stevens. The regression analysis of causal paths. Biometrics, 15(??):236–258, 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Roy:1959:CTR

[91] J. Roy. A class of two replicate incomplete block designs. Biometrics, 15 (??):259–269, 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Geisser:1959:MTT

[92] Seymour Geisser. A method for testing treatment effects in the presence of learning. Biometrics, 15(??):389–395, 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gjeddebaek:1959:CSG

[93] N. F. Gjeddebaek. Contribution to the study of grouped observations. IV. Some comments on simple estimates. Biometrics, 15(??):433–439, 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Robertson:1959:SVG

[94] Alan Robertson. The sampling variance of the genetic correlation coefficient. Biometrics, 15(??):469–485, 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mode:1959:PGV

[95] C. J. Mode and H. F. Robinson. Pleiotropism and the genetic variance and covariance. Biometrics, 15(??):518–537, 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Graybill:1959:CUE

[96] Franklin A. Graybill and R. B. Deal. Combining unbiased estimators. Biometrics, 15(??):543–550, 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Steel:1959:MCR

[97] Robert G. D. Steel. A multiple comparison rank sum test: treatments versus control. Biometrics, 15(??):560–572, 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hartley:1959:SCD

[98] H. O. Hartley. Smallest composite designs for quadratic response surfaces. Biometrics, 15(??):611–624, 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tukey:1959:LPM

[99] John W. Tukey. Little pieces of mixed factorials (abstract 627). Biometrics, 15(??):641–642, ???? 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sandvik:1978:EIM

[100] Leiv Sandvik. Erratum: “Inverse, multiple and sequential sample censuses” (Biometrics 8 (1952), 286–306) by D. G. Chapman. Biometrics, 34(3):523, 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [21].